OpenStudy (anonymous):

if st= 15 and Rt = 40 rs = ?

OpenStudy (e.mccormick):

What part of this is giving you trouble?

OpenStudy (anonymous):

i mo not sure about my answer

OpenStudy (e.mccormick):

What was your answer and how did you get to it?

OpenStudy (anonymous):

my answer is 55 i do it like this . rs =st+rt = 15+40=55 is dat good ?

OpenStudy (e.mccormick):

hmmm... how did you go from st+rt to 15+40?

OpenStudy (anonymous):

cause in my paper it says st= 15 nd rt = 40

OpenStudy (anonymous):

i think its 24

OpenStudy (anonymous):

how ?

OpenStudy (anonymous):

what numbers multiply makes 15 and 40?

OpenStudy (e.mccormick):

Sorry... grabbed the wrong chunk. I meant, "rs =st+rt" how did you make that jump?

OpenStudy (anonymous):

i mean the multiples; its 3x5=15 1x15 the only factor that can go into 40 and 15 is 5

OpenStudy (anonymous):

so 40/5=8

OpenStudy (anonymous):

there we got t=5 s=3

OpenStudy (anonymous):

and R=8

OpenStudy (anonymous):

so Rs= 8x3=24

OpenStudy (e.mccormick):

If the only possibilities are integers, yes.

OpenStudy (anonymous):

salikin is it a multiple choice question? if it is check if 24 is in the choice :o

OpenStudy (anonymous):

nop i had to solve it

OpenStudy (anonymous):

oh.....

OpenStudy (anonymous):

u sure about ur answer

OpenStudy (anonymous):

im pretty sure but i could be wrong :)

OpenStudy (e.mccormick):

For integer answers, that one is correct.

OpenStudy (anonymous):

iight i ll took u ar as right ans

OpenStudy (anonymous):

im pretty sure cuz if fractions and decimals are included then r,s,and t could be many answers not just one answer

OpenStudy (e.mccormick):

Yah, with 3 unknowns in 2 equations you get a parametric answer.

OpenStudy (e.mccormick):

There are also negatives. -5, -3, and -8 but you get the same multiples if they are all negative. That was why I said it was true for integers.